40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11685     Entry time: Tue Oct 13 05:48:39 2015
Author: ericq 
Type: Update 
Category: LSC 
Subject: :/ 

[ericq, Gautam]

Despite our best efforts, the grappa remains out of reach: the DRFPMI was not locked tonight. 

We spent a fair amount of time with the AUX X laser, as it was glitching madly again.

DRMI was finicky until I found some more reliable triggering settings; namely aquiring with AS110Q, but after that transitioning the trigger to the same POP22+POPDC combo as PRCL and MICH. With this in place, the DRMI lock seems really indefinite no matter what CARM seems to do; or at least, I always lost lock due to CARM shenanigans after this. 

The most frustrating part was the fact that I just couldn't cross over the AO path stably. It never "clicked" into high circulating power as it normally does (either in PRFPMI, or how it was last week). Various crossover filters and tweaks were attempted to no avail. Morning traffic starts soon, so we're calling it a night. 

ELOG V3.1.3-